OpenStudy (anonymous):

Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1, 3) m = 2

OpenStudy (anonymous):

we can use the point intercept form here of y-y1 = m (x-x1). (x1, y1) so y-3=2(x-1) solve for y.

OpenStudy (the_fizicx99):

You have your slope so all you have to do is plug in your given values into the point slope formula: Y - y1 = M(x-x1) y -3=2(x-1) Then from there you just have to simplify and solve

OpenStudy (anonymous):

maybe she fell asleep. good luck

OpenStudy (the_fizicx99):

>.<